Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability

Expert replies
by CITI29 » Fri Jul 25, 2008 3:18 pm
If 2 different represenatives are to be selected at random from a group of 10 employee and if P is the probability that both representatives selected will be women, is P >1/2?

1)More than 1/2 of the employees are women.
2)The probability thta both representatives will be men is less than 1/10.

pls someone help.
Join the discussion
Source: — Data Sufficiency |

by reachac » Fri Jul 25, 2008 9:49 pm
Is the ans E?
Join the discussion

by pepeprepa » Sat Jul 26, 2008 12:27 am
I say B

If 2 different represenatives are to be selected at random from a group of 10 employees and if P is the probability that both representatives selected will be women, is P >1/2?

1)More than 1/2 of the employees are women.
Out of 10 employees there are at least 6 women. We cannot assume there are more. So let's see the probability with 6 women.
P=6/10*5/9=1/3 <1/2 Not engouh information
2)The probability that both representatives will be men is less than 1/10.
The maximum number of men is 3 because 3/10*2/9=1/15 and 4/10*3/9=2/15>1/10
P=7/10*6/9=7/15 > 1/2

B
Join the discussion

by CITI29 » Sat Jul 26, 2008 5:05 am
Yes, ans is 'E'...can u pls explain ?
Join the discussion

by cubicle_bound_misfit » Sat Jul 26, 2008 6:01 am
I am getting an answer E

Let us see what makes the event space

(two selected are males) or (two selected are females) or (one selected male and one selected female)

stmt 1 : there can be less than half for 7 women in the group or more than 1/2 for 8 women in the group -------- insufficient.

stmt 2 the probabilty that both representative will be female or one male one female > 9/10 again we do not have a way to solve this case.

Together still we do not have a clue to solve what is the prob if one female-one male is chosen

let me know if my approach is wrong.

regards,
Cubicle Bound Misfit
Join the discussion

by ohwell » Wed Nov 12, 2008 8:00 pm
The answer is E, also because I have the answer.

I followed the same route as pepeprepa, however, pepeprepa's conclusion is not correct. First of all because 7/15 is not > 1/2 but less than 1/2.

I actually think that the maximum number of men can be three and thus there would be seven women. But there could also be just two men (or maybe even just one), then there are 8 women. This can still be a scenario in the second case, it seems to me. In that case:

With 3 men and 7 women, we get 7/10*6/9 < 1/2
With 2 men and 8 women, we get 8/10*7/9 > 1/2

So that is why I would choose E.
Join the discussion